18. How much momentum will a dumb-bell of mass 10 kg transfer to the floor if it falls from a height of 80 cm? Take its downward acceleration to be 10 ms–2.

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
Mass (m) = 10 kg Final velocity (u) = 0 ms–1 Distance (s) = 80 cm = 0.8 m Acceleration (a) = 10 ms–1 By third equation of motion : v2 – u2 = 2as v2 – (0 ms–1)2 = 2 × 10 ms–2 × 0.8 m v = 4 ms–1 Momentum transferred to the floor : = m × v = 10 kg × 4 ms–1 = 40 kg ms–1